Question 4, chap 133, sect 7. part 1 of 1 10 points In a series RLC ac circuit, the resistance is 2811, the inductance is 31 mH, nnd the ca- pacitance is 7 μF. The maximum potential is 117 V, and the angular frequency is 100 rad/s. Caleulate the maximum current in the cir- cuit 1.0.0820623 A 2. 0.0855812 A 3. 0.0882837 A 4. 0.0913059 A 5. 0.0942847 A
Answer #1

Inductive reactance

XL = WL =100*(31*10-3) =3.1 ohms

capacitive reactance

XC=1/WC =1/100*(7*10-6) =1428.57 ohms

Impedance

Z=sqrt[R2+(XL-XC)2]

Z=sqrt[282+(3.1-1428.57)2]

Z=1425.746 ohms

Maximum current in the circuit

I=V/R =117/1425.746

I=0.0820623 A